OpenStudy (anonymous):

CAN I PLZ PLZ GET HELP PLZ I WILL FAN U AND A MEDAL Find the ratio of a to b. 9a = 2b 9 to 2 2 to 11 2 to 9

OpenStudy (anonymous):

i think it's 9 to 2

OpenStudy (anonymous):

I am not sure what they really want, but I assume they are trying to match the two, so if that is the case then it should be 2 to 9 Because well 9*2 18 and 2*9 is 18

OpenStudy (anonymous):

cuz he said a to b

OpenStudy (anonymous):

Yes but they are not equal like the problem wants @meryamium

OpenStudy (anonymous):

thxs guys can u help with more plz

OpenStudy (anonymous):

Yes, but you will have to make a new question because a moderator can come and take you're account down because its a violation to terms of service.

OpenStudy (anonymous):

To have two questions in one actual place

OpenStudy (anonymous):

@samman8801 May be you're right .. can U HELP ME IN MY PROJECT IT'S ALL ABOUT MATH , BUT I DIDN'T UNDERSTAND IT.

OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, 1) and (–2, 5)? oh okay

OpenStudy (anonymous):

I actual have to leave now and attend to my own work :(, if you are still on later and still need help I would be glad to help

OpenStudy (anonymous):

actually**

OpenStudy (anonymous):

okay but thxs for staying @meryamium

OpenStudy (anonymous):

@jazzy288a it's easy \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

and then replace the numbers

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so what would it be A

OpenStudy (er.mohd.amir):

9a=2b a/b = 2/9